OpenStudy (anonymous):

find all the roots if x^4-6x^3+7x^2+6x+1=0

OpenStudy (mathmale):

Mind starting off by sharing your thoughts about approaching this problem? Which methods have you used in the past to find the roots of given polynomial equations?

OpenStudy (mathmale):

Have you seen and used either synthetic division or long division in the past?

OpenStudy (mathmale):

@tom123cruise: your thoughts?

OpenStudy (anonymous):

yeah i tried with the synthetic division... with nos 1 to 4 both -&+ but the remainder is not 0.

OpenStudy (mathmale):

I have to agree that neither -1 nor 1 appears to be a root, as the remainder after synthetic division is not zero in either case. Have you learned any other methods of solving polynomial equations?

OpenStudy (mathmale):

Have you ever heard of Newton's Method?

OpenStudy (mathmale):

As a last resort: Please double check to ensure that you've copied this problem down correctly. Using Newton's Method, I'm getting closer and closer to a root, but that root is a negative fraction.

OpenStudy (anonymous):

sry i dont know newtons method.

OpenStudy (mathmale):

I'd suggest you 1) double check to ensure that you have copied down this problem correctly, and 2) if you don't know Newton's Method, and synthetic division does not succeed, move on to the next problem you have to solve.

OpenStudy (anonymous):

yeah the problem is correctly copied. anyway thanks.

OpenStudy (anonymous):

By the rational root test, the only possible rational solutions are x = 1 or -1, but none of those work.

OpenStudy (anonymous):

I've got it! I think. Using substitution.

OpenStudy (anonymous):

can u please post it here

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

I noticed that the coefficients are almost symmetric if written sequentially: 1, -6, 7, 6, 1. That was the clue that substitution could be used. \(x^4-6x^3+7x^2+6x+1=0\) Let \(x = iy\). \(y^4+6iy^3-7y^2+6iy+1=0\) \(y^2+6iy-7+\frac{6i}y+\frac1{y^2}=0\) Let \(z = y + \frac1y\) and find \(a, b, c\) such that \(az^2+bz+c = y^2+6iy-7+\frac{6i}y+\frac1{y^2}\). \(a = 1,\ b = 6i,\ c = -3\) \(z^2 + 6iz -3 = 0\) Solve for \(z\), substitute back into the relation between \(z\) and \(y\), solve for \(y\), substitute back into the relation between \(y\) and \(x\), solve for \(x\). Really tough. The algebra will take long. Was my explanation clear enough?

OpenStudy (anonymous):

I have to go to bed soon.

OpenStudy (anonymous):

it was gud enough. thanks. im not being rude do u know completing the square method?

OpenStudy (anonymous):

Yes, why?

OpenStudy (anonymous):

can that method can be applied here in this eqn?

OpenStudy (anonymous):

Do you mean the one involving \(z\)?

OpenStudy (anonymous):

wat time is at ur location?? here its pretty late so gotta go. bye and thanks

OpenStudy (anonymous):

Midnight. Goodbye, you're welcome :)

OpenStudy (anonymous):

From Mathematica 9:\[x^4-6 x^3+7 x^2+6 x+1=\left(x^2-3 x-1\right)^2 \]

OpenStudy (anonymous):

5 minutes ago! I just found out too.

OpenStudy (anonymous):

I got the clue that the roots were double roots by looking at the graph and seeing that it was tangential to the x axis at both roots. This meant that we should (maybe) be able to write the polynomial as \(\begin{split} p(x) &= x^4-6x^3+7x^2+6x+1 \\ &= y^4+6y^3-7y^2+6y+1 \\ &= (y^2+ay+1)(y^2+by+1)) \end{split}\) By expanding and comparing coefficients we find that \(a = b = 3i\). Substituting \(x=yi\) we get robtobey's expression.

OpenStudy (anonymous):

Maybe that's a bit too much guesswork though.

OpenStudy (anonymous):

Sorry if you didn't understand my last post. I don't understand my own words. I'll try again. Observe that the polynomial \(p(x) = x^4-6x^3+7x^2+6x+1\) seems to have a local minimum at each of its roots by graphing it. Assume that that's true. Then the multiplicity of each root is even. Because the sum of the multiplicities is the degree of the polynomial (four), the multiplicity of each root is two and we should be able to write \(p(x) = k(x-x_1)^2(x-x_2)^2 = k((x-x_1)(x-x_2))^2 = k(x^2+ax+b)^2\) for some constants \(k, a, b\). Expand \(p(x) = kx^4+2kax^3 + k(a^2+2b)x^2+2kabx+kb^2\). Compare the coefficients of the original expression with the new expression to deduce \(k=1\), \(a=-3\), \(b=-1\).

OpenStudy (anonymous):

You can write the given polynomial as a "polynomial of polynomials": \[\begin{align*}x^4-6x^3+7x^2+6x-1&=(x^4-6x^3+9x^2)+(-2x^2+6x)-1\\ &=(x^2-3x)^2-2(x^2-3x)-1\\ &=((x^2-3x)^2-1)^2 \end{align*}\] like @robtobey said.

OpenStudy (anonymous):

@SithsAndGiggles how did you know to go from the first expression to the second?

OpenStudy (anonymous):

The first three terms kind of looked like a squared binomial, so I rewrote \(7x^2\) as \((9-2)x^2\). The extra \(-2x^2\) happens to let you factor again, and quite nicely.

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

(but the first three -1's should be +1 and the second ^2 on the last line should not be at all)

OpenStudy (anonymous):

Basically it's a matter of completing the square. And yes, sorry, minor typo :)

OpenStudy (anonymous):

So the two highest degree terms determined the "inner polynomial" alone. It doesn't seem very general.

OpenStudy (anonymous):

But efficient, yes.

OpenStudy (anonymous):

Yeah I don't expect this method to work for all \(n\)th order polynomials :/ Life would a lot easier.

OpenStudy (anonymous):

Thanks for the discussion!